The length of a rectangle is 8 feet more than its width. The perimeter of the rectangle is 36 feet. What is the width of the rectangle?
Answer by TimothyLamb(4379)

L = w + 8
2L + 2w = 36
---
2L + 2w = 36
2(w + 8) + 2w = 36
2w + 16 + 2w = 36
4w = 36 - 16
4w = 20
w = 5
L = w + 8
L = 13
---
answer:
w = 5 ft
L = 13 ft
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php